# Watt-class injection-locked diode laser system at 399 nm for atomic physics

### v1

@@ I Introduction
High-power, narrow-linewidth lasers are critical for atomic physics experiments. They enable laser cooling and magneto-optical trapping of atoms and are necessary for precision measurements Kim et al. (2025); Hassan et al. (2025); Morel et al. (2020); Parker et al. (2018); Asenbaum et al. (2020); Kim et al. (2020), quantum simulations Kaufman et al. (2016); Bernien et al. (2017); TAKAHASHI (2022), and quantum information science Deist et al. (2022); Lis et al. (2023). It is challenging to achieve high-power lasers with narrow linewidths at blue or near-ultraviolet (UV) wavelengths. Existing technologies at these wavelengths include frequency-doubled lasers Tinsley et al. (2021), solid-state lasers DeRose et al. (2023), injection-locked single-mode lasers Saxberg et al. (2016), and injection-locked multimode lasers Pagett et al. (2016). Near-UV lasers are becoming increasingly relevant for experiments that utilize alkaline-earth atoms Zhou et al. (2024); Abe et al. (2021) or other species with strong UV transitions Vayninger et al. (2025).
@@
In this work, we use a 5.5mW5.5\ $\text${mW} seed laser to injection lock 57(1)57(1)% of the power of a 1.2W1.2\ $\text${W} “follower” multimode diode laser. The locked power is emitted at the frequency of the seed laser, here tuned to be the $^{1}S_{0}\rightarrow{}^{1}P_{1}$ (398.9nm398.9\ $\text${nm}) transition of ytterbium, with a linewidth only 3.94(6)kHz3.94(6)\ $\text${kHz} broader than that of the seed. The remaining follower power is contained in a broad background spanning 2 THz. We verify these results using heterodyne interferometry and direct atomic spectroscopy of a beam of cold ytterbium.
This system is at least 4 times more powerful than previously demonstrated injection-locked laser systems at 399 nm that are based on single-mode diodes GochnauerThesis . Compared to alternatives that provide similar output power at near-UV wavelengths, the system is notable for its low cost and the ease of availability of its components. Since similar Watt-class multimode diodes are available at other wavelengths, analogous systems can likely be constructed at many wavelengths in the visible and near-UV spectrum.
II System configuration
@@
Figure 1: Schematic design of the injection-locked laser system. Faraday isolators (A) and (B) prevent the follower laser from destabilizing the seed laser and prevent back-reflections from the system output from destabilizing the follower laser. An initial aspheric lens (C) with focal length 13.9mm13.9\ $\text${mm} collects the highly divergent output of the follower diode. Telescopes (D) and (E) circularize and collimate the beam. A scanning Fabry-Perot interferometer (FPI) and laser spectrum analyzer (LSA) monitor the follower beam and provide feedback to the current controller through a computer program. To stabilize the injection lock, a water-cooled manifold maintains the follower laser at 20C20\ $\text${C}.
@@
Figure 2: (A) Normalized power spectra of the seed laser (blue curve) and follower laser (orange, green, and purple curves) measured by the LSA at several follower currents. The unlocked follower spectrum (purple curve) is much broader (2THz2\ $\text${THz}) than shown. (B) Transmission of the follower laser through the FPI with follower current 1.2 A (blue curve) and 0.45 A (orange curve). (C) Normalized spatial intensity profiles of the follower laser at the system output with 39 mW output power (left) and 1000 mW output power (right). Because the follower is a multimode laser, its profile significantly widens and changes shape as the current is increased.
@@
A multimode diode laser (Nichia NDV7975 in a Thorlabs LDM90 mount) serves as the follower. The beam emitted at the diode facet has a non-Gaussian spatial profile with divergence angles of $11$ degrees and $34$ degrees. This highly divergent beam is collected by an aspheric lens, and its astigmatism is then removed by a cylindrical telescope. The beam is then collimated by a circular telescope and passes through a Faraday isolator, which enables its overlap with the seed light and prevents back-reflections from the system output from reaching the follower diode. A small fraction of the follower power is then split into a monitoring branch, where it is sent to a scanning Fabry-Perot interferometer (FPI) and a laser spectrum analyzer (LSA). The signals from these devices are provided to a computer program that adjusts the current controller (Thorlabs LDC4005) supplying the follower laser. We use a Thorlabs SA200-3B with a free spectral range of 1.5 GHz and finesse of 150 as the FPI and a Moglabs Compact Wavemeter as the LSA. The remaining follower beam power is sent to the system output.
@@ II.1 Stabilization and feedback
An injection lock is achieved when the seed laser’s frequency is within the lasing bandwidth of the follower diode and near a resonance of its cavity Lau et al. (2009). The manufacturing process of the follower’s Nichia NDV7975 diode leaves large variation in these diode-medium properties across different units. At ambient conditions of 25C25\ $\text${C}, the diode used here possesses an unlocked spectrum centered at 402nm402\ $\text${nm}. To maintain a lock at the primary 398.9nm398.9\ $\text${nm} ($^{1}S_{0}\rightarrow{}^{1}P_{1}$) transition and to improve stability, a water-cooled manifold was attached to the back of the diode mount to keep the diode at 20C20\ $\text${C}. At this temperature, the central wavelength of the unlocked follower laser is shifted to 400 nm, reducing its frequency difference with the seed laser.
@@
Once established, the lock typically ceases to function within a few minutes without active feedback because of ambient temperature drifts Liu and Slavík (2020); Saxberg et al. (2016), so an active feedback mechanism Milanovic et al. (2025); Chen et al. (2021); Bordonalli et al. (1999) is necessary to achieve a long-term lock. Tuning the follower diode current changes the microscopic properties of the lasing medium, including its length, allowing the follower diode to be kept on resonance with the seed laser Saxberg et al. (2016). A computer program continuously makes adjustments to the follower current to optimize the lock with respect to feedback metrics provided by the LSA and the FPI Saxberg et al. (2016).
@@
Coarse feedback is provided by the LSA (Fig. 2(A)) which measures laser spectra across a 56THz56\ $\text${THz} range with a resolution of 3.8GHz3.8\ $\text${GHz}. As a metric of lock quality, we use the $L^{2}$ inner product of the normalized follower spectrum with the previously recorded seed spectrum. Since the linewidths of both the seed and locked follower are three orders of magnitude smaller than the resolution of the LSA, we find that the LSA alone does not provide a sufficiently precise feedback metric.
@@
Fine feedback is provided by the FPI (Fig. 2(B)). Scanning the FPI’s cavity length changes the frequency classes of the follower spectrum that can transmit through the cavity. When the unlocked follower laser is sent into the cavity, there is no change in transmitted power over the course of the scan because the spectrum has no features finer than the 1.5GHz1.5\ $\text${GHz} free spectral range (see Fig. 2(A), purple curve). When the follower is locked, a significant fraction of its total power is localized near the frequency of the seed laser. Therefore, a transmission peak appears periodically as the cavity length is scanned. The maximum power transmitted over the course of the scan serves as the second feedback metric Saxberg et al. (2016).
@@
To avoid locking to a suboptimal local maximum, the feedback program initially varies the follower current across a broad range and looks for the setpoint that maximizes the $L^{2}$ inner product with the seed spectrum, as measured by the LSA. The program then switches to optimizing the maximum power transmitted through the FPI using gradient descent. In this way, the injection lock has been realized and maintained under ambient temperature fluctuations of up to 2C2\ $\text${C} for over 24 hours.
III Performance
@@
Linewidth. Since the seed and follower spectra are narrower than the 10 MHz resolution of the FPI, a more precise technique is needed to determine their linewidths. For this measurement, we perform heterodyne interferometry with a narrow-linewidth reference laser (Precilaser FL-SF-399-1.5-CW) and measure the interference signal with a 1GHz1\ $\text${GHz} bandwidth photodiode. The reference laser is much narrower than the seed and follower lasers, so the heterodyne response is a direct characterization of their frequency spectra. The results of these measurements are shown in Fig. 3(A, B). The locked follower spectrum is qualitatively similar to the seed spectrum but possesses an additional broad noise background attributable to the unlocked power fraction. This broad background, which can be filtered if necessary Fasano et al. (2021), is subtracted for the purpose of characterizing the spectrum of the locked power.
To characterize differential broadening mechanisms between the follower and seed lasers, we perform heterodyne interferometry between them (Fig. 3(C)). The coherence linewidth between the follower and seed lasers is $\Gamma_{c}=3.94(6)\ \text{kHz}$, indicating that essentially all of the spectral width of the locked follower power is caused by phase noise inherited from the seed laser.
@@
Figure 4: Transmission spectroscopy of a cold beam of ytterbium with (A) the seed and (B) the follower lasers. The yellow curves are pseudo-Voigt fits with Gaussian widths and amplitudes shown. The Lorentzian widths are fixed to the literature value of 29MHz29\ $\text${MHz} Kroeze et al. (2025). The resonances are shifted off their literature values in a manner consistent with a small Doppler shift arising from non-orthogonality between the lasers and atomic beam. Adjusted $R^{2}$ values for each fit are also shown.
To measure the locked power fraction, we perform unsaturated transmission spectroscopy of an ytterbium atomic beam (AOSense Cold Atomic Beam Source) (Fig. 4). As the linewidth of the primary $^{1}S_{0}\rightarrow{}^{1}P_{1}$ transition is 29MHz29\ $\text${MHz} Kroeze et al. (2025) and much narrower than the unlocked power spectrum of the follower laser, the depth of the absorption feature for spectroscopy done with the follower laser is a proxy for the locked power fraction. From the ratio of optical depths measured with the seed and follower lasers, we calculate the locked power fraction to be 0.57(1), which implies a maximum locked power of 680(10)mW680(10)\ $\text${mW} at the diode facet and 550(10)mW550(10)\ $\text${mW} at the system output. Coarser measures of the locked power fraction include the relative strengths of FPI transmission, the power difference measured by the LSA, and the relative amplitudes of the heterodyne resonances with the reference laser. We find each of these techniques to be in broad agreement.
@@
IV Conclusion
We have demonstrated a 398.9nm398.9\ $\text${nm} injection-locked multimode diode laser system with 550(10)mW550(10)\ $\text${mW} maximum locked system output power and 3.94(6)kHz3.94(6)\ $\text${kHz} additional broadening compared to its seed laser. The remaining follower power at the system output is present as a broad (2THz2\ $\text${THz}) background. The system can be used to perform cold atomic beam spectroscopy and would be useful for laser cooling and magneto-optical trapping.
Such a system may find use in future experiments with neutral ytterbium, including quantum simulations TAKAHASHI (2022), tests of the equivalence principle Hartwig et al. (2015), and other searches for new physics Abe et al. (2021); Zhou et al. (2024). In addition, we anticipate that analogous systems could be realized at other wavelengths in the visible and near-UV spectrum. Similar Watt-class multimode diode lasers are available at many wavelengths between $375$ nm and $539$ nm.
V Acknowledgments
We acknowledge funding from the National Science Foundation. M.P. acknowledges support from a William H. Miller III Graduate Fellowship.

### v2

@@ Introduction
High-power, narrow-linewidth lasers are critical for atomic physics experiments. They enable laser cooling and magneto-optical trapping of atoms and are necessary for precision measurements [2, 3, 4, 5, 6, 7], quantum simulations [8, 9, 10], and quantum information science [11, 12]. It is challenging to achieve high-power lasers with narrow linewidths at blue or near-ultraviolet (UV) wavelengths. Existing technologies at these wavelengths include frequency-doubled lasers [13], solid-state lasers [14], injection-locked single-mode lasers [15, 16, 17], and injection-locked multimode lasers [18]. Near-UV lasers are becoming increasingly relevant for experiments that utilize alkaline-earth atoms [19, 20] or other species with strong UV transitions [21].
@@
In this work, we use a 5.5mW5.5\ $\text${mW} seed laser to injection lock 57(1)57(1)% of the power of a 1.2W1.2\ $\text${W} “follower” multimode diode laser. The locked power is emitted at the frequency of the seed laser, here tuned to be the $^{1}S_{0}\rightarrow{}^{1}P_{1}$ (398.9nm398.9\ $\text${nm}) transition of ytterbium, with a linewidth only 3.94(6)kHz3.94(6)\ $\text${kHz} broader than that of the seed. The remaining follower power is contained in a broad background spanning 1.3 THz. We verify these results using heterodyne interferometry and direct atomic spectroscopy of a beam of cold ytterbium.
@@
Figure 1: Schematic design of the injection-locked laser system. Faraday isolators (A) and (B) (transmissivity 0.8, isolation efficiency $30\;\text{dB}$) prevent the follower laser from destabilizing the seed laser and prevent back-reflections from the system output from destabilizing the follower laser. An initial aspheric lens (C) with focal length 13.9mm13.9\ $\text${mm} collects the highly divergent output of the follower diode. Telescopes (D) and (E) circularize and collimate the beam. The beam splitter before the system output sends $5\%$ of the power to a scanning Fabry-Perot interferometer (FPI) and laser spectrum analyzer (LSA), which monitor the follower beam and provide feedback to the current controller through a computer program. To stabilize the injection lock, a water-cooled manifold maintains the follower laser at 20°C20\ $\degree\text${C}.
This system is at least 3 times more powerful than previously demonstrated injection-locked laser systems at 399 nm that are based on single-mode diodes [22, 16] and can be utilized in applications, such as laser cooling and trapping, where single-mode operation is not essential. Compared to alternatives that provide similar output power at near-UV wavelengths, the system is notable for its low cost and the ease of availability of its components. Since similar Watt-class multimode diodes are available at other wavelengths, analogous systems can likely be constructed at many wavelengths in the visible and near-UV spectrum.
@@
Figure 2: (A) Normalized power spectra of the seed laser (blue curve) and follower laser (orange, green, and purple curves) measured by the LSA at several follower currents. The unlocked follower spectrum (purple curve) is much broader ($\Gamma_{\text{FWHM}}=1.3\ \text{THz}$) than the resolution-limited locked follower spectrum ($\Gamma_{\text{FWHM}}=70\ \text{GHz}$). (B) Transmission of the follower laser through the FPI with follower current 1.2 A (blue curve) and 0.45 A (orange curve). (C) Normalized spatial intensity profiles of the follower laser at the system output with 39 mW output power (left) and 1000 mW output power (right). Because the follower is a multimode laser, its profile significantly widens and changes shape as the current is increased.
System configuration
@@
A multimode diode laser (Nichia NDV7975 in a Thorlabs LDM90 mount) serves as the follower. The beam emitted at the diode facet has a non-Gaussian spatial profile with divergence angles of $11$ degrees and $34$ degrees. This highly divergent beam is collected by an aspheric lens, and its astigmatism and ellipticity are then removed by a cylindrical telescope. The beam is then collimated by a circular telescope and passes through a Faraday isolator, which enables its overlap with the seed light and prevents back-reflections from the system output from reaching the follower diode. A small fraction ($5\%$) of the follower power is then split into a monitoring branch, where it is sent to a scanning Fabry-Perot interferometer (FPI) and a laser spectrum analyzer (LSA). The signals from these devices are provided to a computer program that adjusts the current controller (Thorlabs LDC4005) supplying the follower laser. We use a Thorlabs SA200-3B with a free spectral range of 1.5 GHz and finesse of 150 as the FPI and a Moglabs Compact Wavemeter as the LSA. The remaining follower beam power is sent to the system output.
@@
An injection lock is achieved when the seed laser’s frequency is within the lasing bandwidth of the follower diode and near a resonance of its cavity [23]. The manufacturing process of the follower’s Nichia NDV7975 diode leaves large variation in these diode-medium properties across different units. At ambient conditions of 25°C25\ $\degree\text${C}, the diode used here possesses an unlocked spectrum centered at 402nm402\ $\text${nm}. To maintain a lock at the primary 398.9nm398.9\ $\text${nm} ($^{1}S_{0}\rightarrow{}^{1}P_{1}$) transition and to improve stability, a water-cooled manifold was attached to the back of the diode mount to keep the diode at 20°C20\ $\degree\text${C}. At this temperature, the central wavelength of the unlocked follower laser is shifted to 400 nm, reducing its frequency difference with the seed laser.
@@
Once established, the lock typically ceases to function within a few minutes without active feedback because of ambient temperature drifts [24, 15], so an active feedback mechanism [25, 26, 27] is necessary to achieve a long-term lock. Tuning the follower diode current changes the microscopic properties of the lasing medium, including its length, allowing the follower diode to be kept on resonance with the seed laser [15]. The active feedback then consists of a computer program continuously making adjustments to the follower current to optimize the lock with respect to feedback metrics provided by the LSA and the FPI [15].
@@
Coarse feedback is provided by the LSA (Fig. 2(A)) which measures laser spectra across a 56THz56\ $\text${THz} range. As a metric of lock quality, we use the $L^{2}$ inner product of the normalized follower spectrum with the previously recorded seed spectrum. Since the linewidths of both the seed and locked follower are three orders of magnitude smaller than the resolution of the LSA, we find that the LSA alone does not provide a sufficiently precise feedback metric.
@@
Fine feedback is provided by the FPI (Fig. 2(B)). Scanning the FPI’s cavity length changes the frequency classes of the follower spectrum that can transmit through the cavity. When the unlocked follower laser is sent into the cavity, there is no change in transmitted power over the course of the scan because the spectrum has no features finer than the 1.5GHz1.5\ $\text${GHz} free spectral range (see Fig. 2(A), purple curve). When the follower is locked, a significant fraction of its total power is localized near the frequency of the seed laser. Therefore, a transmission peak appears periodically as the cavity length is scanned. The maximum power transmitted over the course of the scan serves as the second feedback metric [15].
@@
To avoid locking to a suboptimal local maximum, the feedback program initially varies the follower current across a broad range and looks for the setpoint that maximizes the $L^{2}$ inner product with the seed spectrum, as measured by the LSA. The program then switches to optimizing the maximum power transmitted through the FPI using gradient descent. In this way, the injection lock has been realized and maintained under ambient temperature fluctuations of up to 2°C2\ $\degree\text${C} for over 24 hours.
Performance
@@
Linewidth. Since the seed and follower spectra are narrower than the 10 MHz resolution of the FPI, a more precise technique is needed to determine their linewidths. For this measurement, we perform heterodyne interferometry with a narrow-linewidth reference laser (Precilaser FL-SF-399-1.5-CW) and measure the interference signal with a 1GHz1\ $\text${GHz} bandwidth photodiode. The reference laser is much narrower than the seed and follower lasers, so the heterodyne response is a direct characterization of their frequency spectra. The results of these measurements are shown in Fig. 3(A, B). The locked follower spectrum is qualitatively similar to the seed spectrum but possesses an additional broad noise background attributable to the unlocked power fraction. This broad background, which can be filtered if necessary [28], is subtracted for the purpose of characterizing the spectrum of the locked power.
To characterize differential broadening mechanisms between the follower and seed lasers, we perform heterodyne interferometry between them (Fig. 3(C)). The coherence linewidth between the follower and seed lasers, given by the inverse of the optical coherence time and calculated as suggested in [29], is $\Gamma_{c}=3.94(6)\ \text{kHz}$, indicating that essentially all of the spectral width of the locked follower power is caused by phase noise inherited from the seed laser.
@@
Figure 4: Transmission spectroscopy of a cold beam of ytterbium with (A) the seed and (B) the follower lasers. The yellow curves are pseudo-Voigt fits with Gaussian widths and amplitudes shown. The Lorentzian widths are fixed to the literature value of 29MHz29\ $\text${MHz} [30]. The resonances are shifted off their literature values in a manner consistent with a small Doppler shift arising from non-orthogonality between the lasers and atomic beam. Adjusted $R^{2}$ values for each fit are also shown. Non-Lorentzian aspects of the line shapes are consistent with uncorrected power drifts in the seed and follower beams.
To measure the locked power fraction, we perform unsaturated transmission spectroscopy of an ytterbium atomic beam (AOSense Cold Atomic Beam Source) (Fig. 4). As the linewidth of the primary $^{1}S_{0}\rightarrow{}^{1}P_{1}$ transition is 29MHz29\ $\text${MHz} [30] and much narrower than the unlocked power spectrum of the follower laser, the depth of the absorption feature for spectroscopy done with the follower laser is a proxy for the locked power fraction. From the ratio of optical depths measured with the seed and follower lasers, we calculate the locked power fraction to be 0.57(1), which implies a maximum locked power of 680(10)mW680(10)\ $\text${mW} at the diode facet and 550(10)mW550(10)\ $\text${mW} at the system output. Coarser measures of the locked power fraction include the relative strengths of FPI transmission, the power difference measured by the LSA, and the relative amplitudes of the heterodyne resonances with the reference laser. We find each of these techniques to be in broad agreement.
@@
The follower has $1\%$ relative intensity noise on a 1s1\ $\text${s} time scale, which is $3.6$ times higher than the seed. This is sufficiently stable for laser cooling and trapping, but for applications where greater intensity control is required, the follower intensity could be actively stabilized.
Conclusion
We have demonstrated a 398.9nm398.9\ $\text${nm} injection-locked multimode diode laser system with 550(10)mW550(10)\ $\text${mW} maximum locked system output power and 3.94(6)kHz3.94(6)\ $\text${kHz} additional broadening compared to its seed laser. The remaining follower power at the system output is present as a broad (1.3THz1.3\ $\text${THz}) background. The system can be used to perform cold atomic beam spectroscopy and would be useful for laser cooling and magneto-optical trapping.
Such a system may find use in future experiments with neutral ytterbium, including quantum simulations [10], tests of the equivalence principle [31], and other searches for new physics [20, 19]. In addition, we anticipate that analogous systems could be realized at other wavelengths in the visible and near-UV spectrum. Similar Watt-class multimode diode lasers are available at many wavelengths between $375$ nm and $539$ nm.
Funding
We acknowledge support from the National Science Foundation (FAIN 2409710). M.P. acknowledges support from a William H. Miller III Graduate Fellowship.
Disclosures
The authors declare no conflicts of interest.
Data availability
Data underlying the results presented in this paper may be obtained from the authors upon reasonable request.